Let the market supply function of 2000 sellers is given as Qs = 4000 + 20Px. and the individual demand function of 5000 buyers is given as 2px = 1500 – 0.2Qd. Then compute

A.  The market clearing price. B.  The equilibrium quantity. C.  Explain graphically.


1
Expert's answer
2022-02-15T08:39:07-0500

At equilibrium: Qd = Qs

"Q_d = 7500 \u2013 10p_x"


"7500 \u2013 10p_x = 4000 + 20p_x"

"7500 \u2013 4000 = 20p_x + 10p_x"

"3500 = 30p_x"

"P_x = 117"

b) Equilibrium quantity

Qd = 7500 – 10(117) = 7500 – 1170 = 6330 units

c) Graphic representation